find an equation for the linear function g(x) which is perpendicular to the line 7x-4y=12 and intersects the line 7x-4y=12 at x=24
please help i can't figure this out and my assignment is due tonight:(

Answer by Alan3354(69443)

find an equation for the linear function g(x) which is perpendicular to the line 7x-4y=12
----------------
Step 1:
Find the slope of the line.
Lines perpendicular have slopes that are negative inverses.
Slope = m
===========
and intersects the line 7x-4y=12 at x=24
--
Step 2:
7x-4y=12 at x=24 --> (24,39)
Use y-y1 = m*(x-x1) where m = slope, and (x1,y1) = (24,39)
--------